OpenStudy (anonymous):

Clarissa works in sales for a software company. Her pay each month varies since each month she receives base pay of $2,400.00 plus 8% commission on her sales. During the month of November, Clarissa had $18,000.00 in sales. For the month of November, ___ % of Clarissa's pay was from her commission.

OpenStudy (doc.brown):

What was her commission in November?

OpenStudy (anonymous):

Idk,18,000.00

OpenStudy (doc.brown):

18000.00 was her sales, you said she made "plus 8% commission on her sales"

OpenStudy (anonymous):

she need to know what as the persentage commission

OpenStudy (doc.brown):

Do you know how to find 8% of 18000.00?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

180%

OpenStudy (doc.brown):

Percent means "of one hundred", so if you chop 18000 into 100 even pieces, Clarissa gets to keep 8 of those pieces.

OpenStudy (anonymous):

18,000-2,400= 15,600 15,600(.08)= $1,248

OpenStudy (doc.brown):

That's not right @TianaChante, Clarissa gets to keep \($2400 +\frac{8}{100}sales\)

OpenStudy (doc.brown):

Or, like I said before @Great_And_Powerful \($2400+(\frac{sales}{100}\times 8)\)